OpenStudy (goformit100):

Find the value of the following cos 1395

OpenStudy (anonymous):

The best way to solve this problem would be a scientific calculator.

OpenStudy (anonymous):

google? I thought "mathletes" were supposed to be better than this.

OpenStudy (anonymous):

\[\cos(1440+45)\]

OpenStudy (shamim):

cos(3*360+315)

OpenStudy (shamim):

cos315

OpenStudy (anonymous):

so did u fine out the answer

OpenStudy (anonymous):

4th quadrant angle

OpenStudy (anonymous):

cos 45 is the answer

OpenStudy (shamim):

cos(360-45)

OpenStudy (shamim):

cos45

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt(2) }\]

OpenStudy (shamim):

@goformit100

OpenStudy (neer2890):

cos(15*90+45) (15*90+45) indicates 4th quadrant. in which cos is +ve. additionally, as the multiple of 90 is odd, so cos changes to sin theta so cos(1395)=sin45

OpenStudy (goformit100):

How please explain ?

OpenStudy (anonymous):

1395 Radians or Degrees?

OpenStudy (neer2890):